# Weak compactness cardinals for strong logics and subtlety properties of the class of ordinals

### v1

@@ 6. Structural properties of abstract logics
We now work towards proving the backward implications in Theorems 1.4, 1.6 and 5.5. Given a natural number $n>0$ and a cardinal $\kappa$ that is not $C^{(n)}$-weakly shrewd, we let $\theta^{n}_{\kappa}$ denote the minimal element $\theta$ of $C^{(n)}$ greater than $\kappa$ with the property that there exists $z\in{\rm{H}}(\theta)$ such that for all $\bar{\theta}\in C^{(n)}$, all cardinals $\bar{\kappa}<\min(\kappa,\bar{\theta})$ and all elementary submodels $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$, there is no elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $z\in{{\rm{ran}}(j)}$. In addition, given a natural number $n>0$ and a cardinal $\kappa$ that is not $C^{(n)}$-weakly shrewd, we let $Z^{n}_{\kappa}$ denote the non-empty set of all elements $z$ of ${\rm{H}}(\theta^{n}_{\kappa})$ with the property that for all $\bar{\theta}\in C^{(n)}$, all cardinals $\bar{\kappa}<\min(\kappa,\bar{\theta})$ and all elementary submodels $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$, there is no elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta^{n}_{\kappa})}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $z\in{{\rm{ran}}(j)}$. In the following, let $\sigma$ denote a language consisting of a binary relation symbol $E$ and four constant symbols $d$, $e$, $f$ and $g$. Given a natural number $m$, a natural number $n>0$ and a cardinal $\rho\in C^{(1)}$ of uncountable cofinality, we define ${\mathcal{C}}^{m,n}_{\rho}$ to be the class of all $\sigma$-structures $M$ with the property that there exists a (necessarily unique) cardinal $\rho<\kappa_{M}\in{\rm{Lim}}(C^{(m)})$ such that the following statements hold:
@@
    Given $M\in Str_{\tau}$ and a first-order sentence $\phi\in\mathcal{L}(\tau)$, then $M\models_{\mathcal{L}}\phi$ holds if and only if $\phi$ holds in $M$ (in the sense of first-order structures).
@@
    Given $M\in Str_{\tau}$, then $M\models_{\mathcal{L}}\phi_{*}$ holds if and only if $M\restriction\sigma$ is isomorphic to a structure in ${\mathcal{C}}^{m,n}_{\rho}$.
@@
    Given $\xi\leq\rho$ and $M\in Str_{\tau}$, then $M\models_{\mathcal{L}}\phi_{{<}\xi}$ holds if and only if $\langle{{\rm{dom}}(M)},E^{M}\rangle$ is a well-order of order-type less than $\xi$.
@@
    Given $\xi\leq\rho$ and $M\in Str_{\tau}$, then $M\models_{\mathcal{L}}\phi_{{\geq}\xi}$ holds if and only if $\langle{{\rm{dom}}(M)},E^{M}\rangle$ is a well-order of order-type at least $\xi$.
@@
Now, assume that $\kappa$ is a strict Löwenheim–Skolem–Tarski number for $\mathcal{L}$. Let $\tau$ be a language extending $\sigma$ that adds functions symbols for first-order Skolem functions to $\sigma$ and let $N$ be $\tau$-structure such that $N\restriction\sigma=M$ and $N$ interprets all new function symbols as Skolem functions. Then $N\models_{\mathcal{L}}\phi_{*}$ and $|{{{\rm{dom}}(N)}}|=\kappa$. Therefore, our assumptions yield a substructure $K$ of $N$ with $K\models_{\mathcal{L}}\phi_{*}$ and $|{{{\rm{dom}}(K)}}|<\kappa$. We can now find a $\sigma$-structure $H$ in ${\mathcal{C}}$ that is isomorphic to $K\restriction\sigma$. Our setup now ensures that there is an elementary embedding ${j}:{{{\rm{dom}}(H)}}\longrightarrow{{{\rm{dom}}(M)}}$ with $j(\kappa_{H})=\kappa$, $j(\rho)=\rho$, $j(f^{H})=f^{M}$ and $j(g^{H})=g^{M}$. Moreover, since $\kappa_{H}=|{{{\rm{dom}}(H)}}|<\kappa$, we know that $j\restriction(\kappa_{H}+1)\neq{\rm{id}}_{\kappa_{H}+1}$. ∎
@@
    If every abstract logic has, for every natural number $n$, either a weak compactness cardinal or a strict Löwenheim–Skolem–Tarski number that is an element of $C^{(n)}$, then ${\rm{Ord}}$ is essentially subtle.
@@
(1) If $n>0$ is a natural number and $\rho$ is a cardinal in $C^{(1)}$ of uncountable cofinality, then there exists an abstract logic $\mathcal{L}$ that satisfies the statements (1)–(5) listed in Lemma 6.4 for $m=0$, $n$ and $\rho$. The fact that $C^{(m)}$ is the class of all ordinals then allows us to apply the lemma to find a $C^{(n)}$-weakly shrewd cardinal greater than $\rho$. By Theorem 3.5, these computations show that ${\rm{Ord}}$ is essentially faint.
@@
In the remainder of this section, we consider the questions whether it is possible to extend the conclusion of Theorem 5.5.(1) to obtain a characterization of the existence of weak Löwenheim–Skolem–Tarski numbers for all abstract logics through the assumption that ${\rm{Ord}}$ is essentially faint.
@@
Our assumption now ensure that there is a weak Löwenheim–Skolem–Tarski number $\kappa$ for $\mathcal{L}$.
@@
Assume, towards a contradiction, that there exists a cardinal $\eta\leq\delta<\kappa$ with $\delta^{\eta}\geq\kappa$. Let $\tau$ denote a language that is obtained from $\sigma$ by first adding a constant symbol $c_{\gamma}$ for each $\gamma<\delta$ and then closing the language under Skolem functions for first order formulas. Pick a sufficiently large cardinal $\theta>\kappa$ and a $\tau$-structure $N$ such that ${{\rm{dom}}(N)}$ is an elementary submodel of ${\rm{H}}(\theta)$ of cardinality $\kappa$ with $\kappa+1\subseteq{{\rm{dom}}(N)}$, $E^{N}={\in}\restriction({{\rm{dom}}(N)}\times{{\rm{dom}}(N)})$, $c_{\gamma}^{N}=\gamma$ for all $\gamma\leq\delta$ and all new function symbols are interpreted as the corresponding Skolem functions. By elementarity, we know that the set ${{\rm{dom}}(N)}\cap{}^{\eta}\delta$ has cardinality $\kappa$. Since $\kappa$ is a weak Löwenheim–Skolem–Tarski number for $\mathcal{L}$, we can find an $\mathcal{L}$-elementary submodel $M$ of $N$ with $|{{{\rm{dom}}(M)}}|<\kappa$. Then $\bar{M}$ is an elementary submodel of ${\rm{H}}(\theta)$ with $\delta+1\subseteq\bar{M}$. Moreover, we know that there is a function $s\in N\cap{}^{\eta}\delta$ that is not an element of $M$. But, the set $\mathcal{L}_{\rho,\omega}(\tau)$ contains a sentence $\phi_{s}$ that states that there is a function from $c_{\eta}$ to $c_{\delta}$ that sends $c_{\gamma}$ to $c_{s(\gamma)}$ for all $\gamma<\eta$. It then follows that $N\models_{\mathcal{L}}\phi_{s}$ holds and $M\models_{\mathcal{L}}\phi_{s}$ fails, contradicting the fact that $M$ is an $\mathcal{L}$-elementary submodel of $N$. ∎
@@
The above result leaves open the question of whether the existence of weak Löwenheim–Skolem–Tarski numbers for all abstract logics is provably equivalent to the assumption that ${\rm{Ord}}$ is essentially faint (see Question 7.1 in the next section). Below, it is shown that a separation of these principles would require substantial large cardinal assumptions. Recall that the Singular Cardinal Hypothesis ${\rm{SCH}}$ states that $\kappa^{{{\rm{cof}}(\kappa)}}=\kappa^{+}$ holds for every singular cardinal $\kappa$ with $2^{{{\rm{cof}}(\kappa)}}<\kappa$. In the following, we say that the ${\rm{SCH}}$ eventually holds if this implication holds for eventually all singular cardinals. Remember that a classical result of Solovay in [20] shows that the existence of a strongly compact cardinal implies that the ${\rm{SCH}}$ eventually holds.
@@
In one direction, a combination of Theorem 3.5 with Theorem 6.6 directly shows that (1) implies (2). In the other direction, note that, if the ${\rm{SCH}}$ eventually holds, then the proof of [10, Theorem 5.22] shows that for every cardinal $\mu$, all sufficiently large weakly inaccessible cardinals $\kappa$ have the property that $\nu^{{<}\mu}<\kappa$ holds for all $\nu<\kappa$. In particular, it follows that all sufficiently large $C^{(n)}$-weakly shrewd cardinals satisfy the assumptions of (6) in Lemma 6.1 and therefore these cardinals are weak Löwenheim–Skolem–Tarski numbers for abstract logic definable by $\Sigma_{n}$-formulas. This shows that (2) implies (1). ∎
@@ 7. Open questions
Note that, as shown in the proof of Corollary 6.7, a negative answer to Question 7.1 implies the consistency of large cardinal assumptions in the region of measurable cardinals of high Mitchell order, because it implies that the ${\rm{SCH}}$ fails at a proper class of singular cardinals and such a failure is known to imply the existence of large cardinals in canonical inner models (see, for example, [18]). However, it is not clear if the given fragment of the ${\rm{SCH}}$ outright follows from the assumption that ${\rm{Ord}}$ is essentially faint.

### v3

@@ 6. Structural properties of abstract logics
We now work towards proving the backward implications in Theorems 1.4, 1.6 and 5.5. Given a natural number $n>0$ and a cardinal $\kappa$ that is not $C^{(n)}$-weakly shrewd, we let $\theta^{n}_{\kappa}$ denote the minimal element $\theta$ of $C^{(n)}$ greater than $\kappa$ with the property that there exists $z\in{\rm{H}}(\theta)$ such that for all $\bar{\theta}\in C^{(n)}$, all cardinals $\bar{\kappa}<\min(\kappa,\bar{\theta})$ and all elementary submodels $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$, there is no elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta)}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $z\in{{\rm{ran}}(j)}$. In addition, given a natural number $n>0$ and a cardinal $\kappa$ that is not $C^{(n)}$-weakly shrewd, we let $Z^{n}_{\kappa}$ denote the non-empty set of all elements $z$ of ${\rm{H}}(\theta^{n}_{\kappa})$ with the property that for all $\bar{\theta}\in C^{(n)}$, all cardinals $\bar{\kappa}<\min(\kappa,\bar{\theta})$ and all elementary submodels $X$ of ${\rm{H}}(\bar{\theta})$ with $\bar{\kappa}+1\subseteq X$, there is no elementary embedding ${j}:{X}\longrightarrow{{\rm{H}}(\theta^{n}_{\kappa})}$ with $j\restriction\bar{\kappa}={\rm{id}}_{\bar{\kappa}}$, $j(\bar{\kappa})=\kappa$ and $z\in{{\rm{ran}}(j)}$. In the following, let $\sigma$ denote the language consisting of a binary relation symbol $E$ and four constant symbols $d$, $e$, $f$ and $g$. Given a natural number $m$, a natural number $n>0$ and a cardinal $\rho\in C^{(1)}$ of uncountable cofinality, we define ${\mathcal{C}}^{m,n}_{\rho}$ to be the class of all $\sigma$-structures $M$ with the property that there exists a (necessarily unique) cardinal $\rho<\kappa_{M}\in{\rm{Lim}}(C^{(m)})$ such that the following statements hold:
@@
    Given $M\in Str_{\tau}$ and a first-order sentence $\phi\in\mathcal{L}(\tau)$, $M\models_{\mathcal{L}}\phi$ holds if and only if $\phi$ holds in $M$ (in the sense of first-order structures).
@@
    Given $M\in Str_{\tau}$, $M\models_{\mathcal{L}}\phi_{*}$ holds if and only if $M\restriction\sigma$ is isomorphic to a structure in ${\mathcal{C}}^{m,n}_{\rho}$.
@@
    Given $\xi\leq\rho$ and $M\in Str_{\tau}$, $M\models_{\mathcal{L}}\phi_{{<}\xi}$ holds if and only if $\langle{{\rm{dom}}(M)},E^{M}\rangle$ is a well-order of order-type less than $\xi$.
@@
    Given $\xi\leq\rho$ and $M\in Str_{\tau}$, $M\models_{\mathcal{L}}\phi_{{\geq}\xi}$ holds if and only if $\langle{{\rm{dom}}(M)},E^{M}\rangle$ is a well-order of order-type at least $\xi$.
@@
Now, assume that $\kappa$ is a strict Löwenheim–Skolem–Tarski number for $\mathcal{L}$. Let $\tau$ be a language extending $\sigma$ that adds functions symbols for first-order Skolem functions to $\sigma$ and let $N$ be a $\tau$-structure such that $N\restriction\sigma=M$ and $N$ interprets all new function symbols as Skolem functions. Then $N\models_{\mathcal{L}}\phi_{*}$ and $|{{{\rm{dom}}(N)}}|=\kappa$. Therefore, our assumptions yield a substructure $K$ of $N$ with $K\models_{\mathcal{L}}\phi_{*}$ and $|{{{\rm{dom}}(K)}}|<\kappa$. We can now find a $\sigma$-structure $H$ in ${\mathcal{C}}^{m,n}_{\rho}$ that is isomorphic to $K\restriction\sigma$. Our setup now ensures that there is an elementary embedding ${j}:{{{\rm{dom}}(H)}}\longrightarrow{{{\rm{dom}}(M)}}$ with $j(\kappa_{H})=\kappa$, $j(\rho)=\rho$, $j(f^{H})=f^{M}$ and $j(g^{H})=g^{M}$. Moreover, since $\kappa_{H}=|{{{\rm{dom}}(H)}}|<\kappa$, we know that $j\restriction(\kappa_{H}+1)\neq{\rm{id}}_{\kappa_{H}+1}$. ∎
@@
    If every abstract logic has, for every natural number $n$, either a weak compactness cardinal that is an element of $C^{(n)}$ or a strict Löwenheim–Skolem–Tarski number that is an element of $C^{(n)}$, then ${\rm{Ord}}$ is essentially subtle.
@@
(1) Fix a natural number $n>0$ and a cardinal $\rho$ in $C^{(1)}$ of uncountable cofinality. Set $m=0$. Then there exists an abstract logic $\mathcal{L}$ that satisfies the statements (1)–(5) listed in Lemma 6.4 for $m$, $n$ and $\rho$. The fact that $C^{(m)}$ is the class of all ordinals now allows us to apply the lemma to find a $C^{(n)}$-weakly shrewd cardinal greater than $\rho$.
These computations show that for every natural number $n>0$, there is a proper class of $C^{(n)}$-weakly shrewd cardinals. In this situation, Theorem 3.5 allows us to conclude that ${\rm{Ord}}$ is essentially faint.
@@
Remark 6.6.
In [8], Fuchino and Sakai consider the possibility of characterizing the existence of weak compactness cardinals for second-order logic $\mathcal{L}^{2}$ through the existence of certain large cardinals, analogous to Magidor’s characterization of the existence of strong compactness cardinals for $\mathcal{L}^{2}$ through the existence of extendible cardinals in [14]. For this purpose, they introduce the concept of weak extendibility as a natural candidate for a large cardinal property providing such a characterization. Here, a cardinal $\kappa$ is defined to be weakly extendible if $2^{{<}\kappa}=\kappa$ holds and for every ordinal $\theta>\kappa$ and every elementary submodel $X$ of ${\rm{V}}_{\theta}$ of cardinality $\kappa$ with $\kappa+1\subseteq X$, there exists an ordinal $\vartheta>\kappa$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{V}}_{\vartheta}}$ with $j\restriction\kappa={\rm{id}}_{\kappa}$ and $j(\kappa)>\kappa$. The results of [8] show that all weakly extendible cardinals are weakly compact and therefore strongly inaccessible.
Fuchino and Sakai prove that a cardinal $\kappa$ is a weak compactness cardinal for $\mathcal{L}^{2}$ if and only if for every ordinal $\theta>\kappa$ and every elementary submodel $X$ of ${\rm{V}}_{\theta}$ of cardinality $\kappa$ with $\kappa+1\subseteq X$, there exists an ordinal $\vartheta>\kappa$ and an elementary embedding ${j}:{X}\longrightarrow{{\rm{V}}_{\vartheta}}$ with $j(\kappa)>\sup(j[\kappa])$. This equivalence directly implies that every weakly extendible cardinal is a weak compactness cardinal for $\mathcal{L}^{2}$. Moreover, a standard application of Jensen’s Covering Lemma (as in [10, Theorem 18.27]) allows us to conclude that the assumption ${\rm{V}}={\rm{L}}$ causes all weak compactness cardinals for $\mathcal{L}^{2}$ to be weakly extendible. In addition, a standard argument shows that, if $\kappa$ is the least weak compactness cardinal for $\mathcal{L}^{2}$ and $2^{{<}\kappa}=\kappa$ holds, then $\kappa$ is weakly extendible. These implications directly raise the question of whether the existence of a weak compactness cardinal for $\mathcal{L}^{2}$ provably implies the existence of a weakly extendible cardinal. A combination of Theorems 1.6 and 1.14 now provides a negative answer to this question by showing that weak compactness cardinals for $\mathcal{L}^{2}$ can exist in models of ${\rm{ZFC}}$ without strongly inaccessible cardinals.
In the remainder of this section, we consider the question whether it is possible to extend the conclusion of Theorem 5.5.(1) to obtain a characterization of the existence of weak Löwenheim–Skolem–Tarski numbers for all abstract logics through the assumption that ${\rm{Ord}}$ is essentially faint.
@@
Our assumption now ensures that there is a weak Löwenheim–Skolem–Tarski number $\kappa$ for $\mathcal{L}$.
@@
Assume, towards a contradiction, that there exists a cardinal $\eta\leq\delta<\kappa$ with $\delta^{\eta}\geq\kappa$. Let $\tau$ denote a language that is obtained from $\sigma$ by first adding a constant symbol $c_{\gamma}$ for each $\gamma\leq\delta$ and then closing the language under Skolem functions for first order formulas. Pick a sufficiently large cardinal $\theta>\kappa$ and a $\tau$-structure $N$ such that ${{\rm{dom}}(N)}$ is an elementary submodel of ${\rm{H}}(\theta)$ of cardinality $\kappa$ with $\kappa+1\subseteq{{\rm{dom}}(N)}$, $E^{N}={\in}\restriction({{\rm{dom}}(N)}\times{{\rm{dom}}(N)})$, $c_{\gamma}^{N}=\gamma$ for all $\gamma\leq\delta$ and all new function symbols are interpreted as the corresponding Skolem functions. By elementarity, we know that the set ${{\rm{dom}}(N)}\cap{}^{\eta}\delta$ has cardinality $\kappa$. Since $\kappa$ is a weak Löwenheim–Skolem–Tarski number for $\mathcal{L}$, we can find an $\mathcal{L}$-elementary submodel $M$ of $N$ with $|{{{\rm{dom}}(M)}}|<\kappa$. Then ${{\rm{dom}}(M)}$ is an elementary submodel of ${\rm{H}}(\theta)$ with $\delta+1\subseteq{{\rm{dom}}(M)}$. Moreover, we know that there is a function $s\in N\cap{}^{\eta}\delta$ that is not an element of $M$. But, the set $\mathcal{L}_{\rho,\omega}(\tau)$ contains a sentence $\phi_{s}$ that states that there is a function from $c_{\eta}$ to $c_{\delta}$ that sends $c_{\gamma}$ to $c_{s(\gamma)}$ for all $\gamma<\eta$. It then follows that $N\models_{\mathcal{L}}\phi_{s}$ holds and $M\models_{\mathcal{L}}\phi_{s}$ fails, contradicting the fact that $M$ is an $\mathcal{L}$-elementary submodel of $N$. ∎
@@
The above result leaves open the question of whether the existence of weak Löwenheim–Skolem–Tarski numbers for all abstract logics is provably equivalent to the assumption that ${\rm{Ord}}$ is essentially faint (see Question 7.1 in the next section). In the following, we will use Theorem 6.7 to severely restrict the class of models of set theory in which the given principles are not equivalent. For this purpose, recall that the Singular Cardinal Hypothesis ${\rm{SCH}}$ states that $\kappa^{{{\rm{cof}}(\kappa)}}=\kappa^{+}$ holds for every singular cardinal $\kappa$ with $2^{{{\rm{cof}}(\kappa)}}<\kappa$. In the following, we say that the ${\rm{SCH}}$ eventually holds if this implication holds for eventually all singular cardinals.
@@
In one direction, a combination of Theorem 3.5 with Theorem 6.7 directly shows that (1) implies (2). In the other direction, note that, if the ${\rm{SCH}}$ eventually holds, then the proof of [10, Theorem 5.22] shows that for every cardinal $\mu$, all sufficiently large weakly inaccessible cardinals $\kappa$ have the property that $\nu^{{<}\mu}<\kappa$ holds for all $\nu<\kappa$. In particular, it follows that all sufficiently large $C^{(n)}$-weakly shrewd cardinals satisfy the assumptions of (6) in Lemma 6.1 and therefore these cardinals are weak Löwenheim–Skolem–Tarski numbers for abstract logics definable by $\Sigma_{n}$-formulas. This shows that (2) implies (1). ∎
Remember that a classical result of Solovay in [20] shows that the existence of a strongly compact cardinal implies that the ${\rm{SCH}}$ eventually holds. In particular, the existence of such a cardinal entails that ${\rm{Ord}}$ is essentially faint if and only if every abstract logic has a weak Löwenheim–Skolem–Tarski number. In another direction, Corollary 6.8 also shows that the inequivalence of the two principles implies the consistency of large cardinal assumptions in the region of measurable cardinals of high Mitchell order. This follows from the fact that this inequivalence requires a failure of the ${\rm{SCH}}$ at a proper class of singular cardinals and such a failure is known to imply the existence of such large cardinals in canonical inner models (see, for example, [18]).
@@ 7. Open questions
As discussed at the end of Section 6, a negative answer to Question 7.1 would require substantial large cardinal assumptions. However, it is not clear if the given fragment of the ${\rm{SCH}}$ outright follows from the assumption that ${\rm{Ord}}$ is essentially faint.